Determine the pressure drop along a venue of the length L and the radius R, if the viscosity of blood is n and the speed of blood at the central layer of the v enule is Vo


1
Expert's answer
2020-09-24T11:10:05-0400

According to Poiseuille’s law:


"Q=\frac{\Delta p \pi R^4}{8\eta L}"

The flow can be roughly expressed in terms of the speed of blood at the central layer:


"Q=\pi R^2v_0"

Therefore, the pressure drop is


"\Delta p=\frac{8\eta Lv_0}{R^2}"